OpenStudy (xxsagexx):

x ⋅ x = ? (1 point) 2x x2 x(x + 2) x + 2

OpenStudy (xxsagexx):

@triciaal @mathstudent55

OpenStudy (gracygirl):

What do you think?

OpenStudy (xxsagexx):

Im really not sure at all, I have no guesses... I just need help figuring it out (not just a plain answer, I need it to e explained)

OpenStudy (jiteshmeghwal9):

\[a^b.a^c=a^{b+c}\]

OpenStudy (jiteshmeghwal9):

\[x^1.x^1=?\]

OpenStudy (xxsagexx):

I dont understand, can you plz explain the text your posting?

OpenStudy (ajspeller):

When multiplying factors with like bases you have to add the exponents.

OpenStudy (xxsagexx):

Oh, well I dont think there are in exponents in this question. Not that I can see.

OpenStudy (xxsagexx):

any*

OpenStudy (jiteshmeghwal9):

Ahh no \[x^1.x^1=x^{1+1}\]

OpenStudy (gracygirl):

Okay the answer is going to be \[x ^{2}\] because... \[x ^{1} * x ^{1} = x ^{2}\] The ones aren't listed but \[x ^{1} = x\]

OpenStudy (gracygirl):

Because it's only 1 x.

OpenStudy (gracygirl):

Make sense?

OpenStudy (xxsagexx):

Yes, thank you Gracy!

OpenStudy (ajspeller):

Missing exponents are implied to be 1 ..

OpenStudy (gracygirl):

Exactly ^^ Anytime :)